Let 𝑛 n be a positive integer that is at least 3 . Show that 𝑈(2^𝑛) has at least three elements of order 2. Prove using Bezout's theorem


1
Expert's answer
2021-03-09T02:02:30-0500

U(m) is the group of positive integers j "\leq" m such that gcd(j, m) = 1, under multiplication modulo m.

Since elements in u("2^n)" are coprime to 2,

"U(2^n)= {{1, 3, 5...2^n-3, 2^n-1}}" }The order

u("2^n)" is "\varphi(2^n)=2^{n-1}, \\"

"2^{n-1}" is of order 2, since

"(2^{n-1})^2=2^{2n}-2^{n+1} \equiv1mod2^n\\
Also\\
(2^{n-1}+1)^2=2^{2n-2}+2^{n}+1\equiv1mod2^n\\
More so\\
(2^{n-1}-1)^2=2^{n-2}-2^{n}+1\equiv1mod2^n\\
for \\
n\ge3"